# Enhanced DACER Algorithm with High Diffusion Efficiency

### v1

@@ Appendix B Environmental Details
This is a high-performance physics simulation platform widely adopted for robotic reinforcement learning research. The environment features efficient physics computation, accurate dynamic system modeling, and comprehensive support for articulated robots, making it an ideal benchmark for RL algorithm development. In this research, we concentrate on eight tasks: Humanoid, Ant, HalfCheetah, Walker2d, InvertedDoublePendulum (IDP), Hopper, Pusher, and Swimmer, as illustrated in Fig. 6. The IDP task entails maintaining the balance of a double pendulum in an inverted state. In contrast, the objective of the other tasks is to maximize the forward velocity while avoiding falling. All these tasks are realized through the OpenAI Gym interface [2].
Figure 6: MuJoCo simulation tasks. (a) Humanoid-v3:$(s\times a)\in\mathbb{R}^{376}\times\mathbb{R}^{17}$. (b) Ant-v3: $(s\times a)\in\mathbb{R}^{111}\times\mathbb{R}^{8}$. (c) HalfCheetah-v3 : $(s\times a)\in\mathbb{R}^{17}\times\mathbb{R}^{6}$. (d) Walker2d-v3: $(s\times a)\in\mathbb{R}^{17}\times\mathbb{R}^{6}$. (e) InvertedDoublePendulum-v3: $(s\times a)\in\mathbb{R}^{6}\times\mathbb{R}^{1}$. (f) Hopper-v3: $(s\times a)\in\mathbb{R}^{11}\times\mathbb{R}^{3}$. (g) Pusher-v2: $(s\times a)\in\mathbb{R}^{23}\times\mathbb{R}^{7}$. (h) Swimmer-v3: $(s\times a)\in\mathbb{R}^{8}\times\mathbb{R}^{2}$.
Appendix C Additional multimodal experiments
The MuJoCo benchmarks also demonstrate multimodal action distributions across robotic control tasks, as shown in Fig. 7. By leveraging the trained DACER2 to conduct multiple action samplings in identical states, it can be observed that the policy manifests multiple behavioral modes in specific states. Each subfigure depicts 100 actions sampled for the respective state, projected into $\mathbb{R}^{2}$ using UMAP [22] for visualization.
We conducted the same experiment with DACER and observed that while multimodal behaviors are still present in some tasks, the Ant and Walker2d task fails to show clear multimodal patterns under the DACER policy. These results suggest that DACER2 exhibits stronger and more consistent multimodality than DACER across MuJoCo tasks.
Table 2: Hyperparameters used in DACER2.
Parameter Humanoid Ant HalfCheetah Walker2d
Loss function weight $\eta$ 2.0 2.0 2.0 1.0
Parameter IDP Hopper Pusher Swimmer
Loss function weight $\eta$ 0.1 0.1 0.1 0.001
@@ Appendix D Experimental Hyperparameters
The hyperparameters of all baseline algorithms except the diffusion-based algorithm are shown in Table 3. Additionally, the parameters for all diffusion-based algorithms, including DACER2, are presented in Table 2 and Table 4.
@@
   Replay buffer capacity 2,000,000
@@
   Actor learning rate $41{\rm{e-}}$ 4
   Critic learning rate $41{\rm{e-}}$ 4
@@
   Learning rate of $\alpha$ $43{\rm{e-}}$ 4
   Expected entropy ($\overline{\mathcal{H}}$) ¯$\overline{\mathcal{H}}=-{\rm{dim}}(\mathcal{A})$
@@
Table 4: Diffusion-based algorithms’ hyperparameter
Parameter DACER2 DACER QVPO QSM DIPO
Replay buffer capacity 2e6 2e6 2e6 2e6 2e6
Buffer warm-up size 3e4 3e4 3e4 3e4 3e4
Batch size 256 256 256 256 256
Discount $\gamma$ 0.99 0.99 0.99 0.99 0.99
Target network soft-update rate $\rho$ 0.005 0.005 0.005 0.005 0.005
Network update times per iteration 1 1 1 1 1
Action bound $[-1,1]$ $[-1,1]$ $[-1,1]$ $[-1,1]$ $[-1,1]$
Reward scale 0.2 0.2 0.2 0.2 0.2
No. of hidden layers 2 2 2 2 2
No. of hidden nodes 256 256 256 256 256
Activations in critic network GeLU GeLU Mish ReLU Mish
Activations in actor network Mish Mish Mish ReLU Mish
Diffusion steps 5 20 20 20 20
Policy delay update 2 2 2 2 2
Action gradient steps N/A N/A N/A N/A 20
No. of Gaussian distributions 3 3 N/A N/A N/A
No. of action samples 200 200 N/A N/A N/A
Time-weighted hyperparameter $c$ 0.4 N/A N/A N/A N/A
Time-weighted hyperparameter $d$1.8 N/A N/A N/A N/A
Alpha delay update 10,000 10,000 N/A N/A N/A
Noise scale $\lambda$ 0.1 0.1 N/A N/A N/A
Optimizer Adam Adam Adam Adam Adam
Actor learning rate $1\cdot 10^{-4}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$
Critic learning rate $1\cdot 10^{-4}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$
Alpha learning rate $3\cdot 10^{-2}$ $3\cdot 10^{-2}$ N/A N/A N/A
Target entropy $-{\rm{dim}}(\mathcal{A})$ $-{\rm{dim}}(\mathcal{A})$ N/A N/A N/A
Figure 7: Demonstration of multimodal action distributions on MuJoCo benchmarks. From top to bottom they are Humanoid-v3, Walker2d-v3, HalfCheetah-v3, Ant-v3. The figure displays sampled actions from DACER2. Each subfigure depicts 100 actions sampled for the respective state, projected into $\mathbb{R}^{2}$ using UMAP [22].
@@ Appendix E Limitation and Future Work
In this study, we propose the Q-gradient field objective as an auxiliary training loss to guide the diffusion policy with more informative gradient signals. Although we employ a normalization technique on $\nabla_{\boldsymbol{a}}Q(\boldsymbol{s},\boldsymbol{a})$ to enhance training stability, the errors and instability inherent in Q-value estimation can still adversely affect the training process. In future work, we plan to incorporate techniques such as ensemble Q-learning and multimodal value function estimation to mitigate Q-value estimation errors and instability, thereby further improving the robustness of our algorithm.

### v2

@@ Appendix C Environmental Details
This is a high-performance physics simulation platform widely adopted for robotic reinforcement learning research. The environment features efficient physics computation, accurate dynamic system modeling, and comprehensive support for articulated robots, making it an ideal benchmark for RL algorithm development. In this research, we concentrate on eight tasks: Humanoid, Ant, HalfCheetah, Walker2d, InvertedDoublePendulum (IDP), Hopper, HumanoidStandup, and Swimmer. The IDP task entails maintaining the balance of a double pendulum in an inverted state. In contrast, the objective of the other tasks is to maximize the forward velocity while avoiding falling. All these tasks are realized through the OpenAI Gym interface (Brockman, 2016).
Figure 5: Walker2d-v3
State-action space: $\mathcal{S}\in\mathbb{R}^{17}$, $\mathcal{A}\in\mathbb{R}^{6}$.
Objective. Maintain forward velocity as fast as possible while avoiding falling over.
Initialization. The walker is initialized in a standing position with slight random noise added to joint positions and velocities.
Termination. The episode ends when the agent falls, the head touches the ground, or after 1000 steps.
Figure 6: Humanoid-v3
State-action space: $\mathcal{S}\in\mathbb{R}^{376}$, $\mathcal{A}\in\mathbb{R}^{17}$.
Objective. Maintain balance and walk or run forward at a high velocity while avoiding falls.
Initialization. The humanoid starts in an upright position with slight random perturbations to joint angles and velocities.
Termination. The episode ends when the head height is less than 1.0 meter, the torso tilts excessively, or after 1000 steps.
Figure 7: Ant-v3
State-action space: $\mathcal{S}\in\mathbb{R}^{111}$, $\mathcal{A}\in\mathbb{R}^{8}$.
Objective. Navigate forward as quickly as possible using four legs while maintaining stability.
Initialization. The ant is initialized in a stable, upright position with random noise applied to its joints.
Termination. The episode ends if the ant falls, flips over, or reaches the maximum step count of 1000.
Figure 8: Halfcheetah-v3
State-action space: $\mathcal{S}\in\mathbb{R}^{17}$, $\mathcal{A}\in\mathbb{R}^{6}$.
Objective. Achieve maximum forward velocity with smooth, coordinated movements.
Initialization. The agent starts with a slight forward tilt and randomized joint noise.
Termination. The episode ends after 1000 steps or if the agent’s head touches the ground.
Figure 9: Swimmer-v3
State-action space: $\mathcal{S}\in\mathbb{R}^{8}$, $\mathcal{A}\in\mathbb{R}^{2}$.
Objective. Propel forward through water-like dynamics using sinusoidal wave patterns.
Initialization. The swimmer starts in a straight posture with minor random perturbations.
Termination. The episode ends after 1000 steps, with no explicit termination for falling.
Figure 10: HumanoidStandup
State-action space: $\mathcal{S}\in\mathbb{R}^{348}$, $\mathcal{A}\in\mathbb{R}^{17}$.
Objective. Stand up from lying on the ground by applying torques to the joints, with rewards for upward movement and penalties for large actions or strong impacts.
Initialization. The humanoid starts lying down, with small random noise added to joint positions and velocities.
Termination. The episode does not terminate early; it ends after 1000 steps.
Figure 11: Hopper-v3
State-action space: $\mathcal{S}\in\mathbb{R}^{11}$, $\mathcal{A}\in\mathbb{R}^{3}$.
Objective. Hop forward as fast as possible by applying torques to the thigh, leg, and foot joints, while staying upright.
Initialization. The hopper starts standing upright with small random perturbations in position and velocity.
Termination. The episode ends if the hopper falls (body hits the ground) or after 1000 steps.
Figure 12: IDP-v3
State-action space: $\mathcal{S}\in\mathbb{R}^{9}$, $\mathcal{A}\in\mathbb{R}^{1}$.
Objective. Balance the second pole upright by applying horizontal forces to the cart, while maximizing time alive and minimizing tip distance and joint velocities.
Initialization. The cart and poles start near the upright position with small random noise in position and velocity.
Termination. The episode ends if the tip of the second pole falls below height 1. Otherwise, it is truncated after 1000 steps.
@@ Appendix D Experimental Hyperparameters
The hyperparameters of all baseline algorithms except the diffusion-based algorithm are shown in Table 4. Additionally, the parameters for all diffusion-based algorithms, including DACERv2, are presented in Table 5 and Table 6.
The hyperparameter $c,d$ for time-weighted mechanism is determined by the diffusion step size, inspired by the variance-preserving beta schedule used in DDPM (Ho et al., 2020). The code of implementation is as follows:
def vp_alpha_schedule(timesteps: int, b_min=0.1, b_max=10.):
    T = timesteps
    t = np.arange(1, T + 1)
    return np.exp(-b_min / T - 0.5 * (b_max - b_min) * (2 * t - 1) / T ** 2)
# Set parameters
timesteps = 5
alphas = vp_alpha_schedule(timesteps)
# Reverse the alpha array as in B.alphas[self.agent.num_timesteps - 1 - t]
reversed_alphas = alphas[::-1]
t_vals = np.arange(timesteps)
# Fit the exponential form exp(ct + d)
params, _ = curve_fit(exp_fit, t_vals, reversed_alphas)
c, d = params
@@
   Replay buffer capacity 1,000,000
@@
   Actor learning rate $1{\rm{e-}}4$
   Critic learning rate $1{\rm{e-}}4$
@@
   Learning rate of $\alpha$ $3{\rm{e-}}4$
   Expected entropy ($\overline{\mathcal{H}}$) $\overline{\mathcal{H}}=-{\rm{dim}}(\mathcal{A})$
@@
Table 5: Hyperparameter $\eta$ used in DACERv2.
Task Hopper Ant HalfCheetah Walker2d MultiGoal Hum. S. Humanoid Swimmer IDP
$\eta$ 1.0 1.0 1.0 1.0 1.0 0.01 0.01 0.01 0.01
Table 6: Diffusion-based algorithms’ hyperparameters
Parameter DACERv2 DACER QVPO QSM DIME DIPO
Replay buffer capacity 1e6 1e6 1e6 1e6 1e6 1e6
Buffer warm-up size 3e4 3e4 3e4 3e4 3e4 3e4
Batch size 256 256 256 256 256 256
Discount $\gamma$ 0.99 0.99 0.99 0.99 0.99 0.99
Target network soft-update rate $\rho$ 0.005 0.005 0.005 0.005 N/A 0.005
Network update times per iteration 1 1 1 1 1 1
Action bound [1,1][-1,1] [1,1][-1,1] [1,1][-1,1] [1,1][-1,1] [1,1][-1,1] [1,1][-1,1]
Reward scale 0.2 0.2 0.2 0.2 0.2 0.2
No. of Actor layers 2 2 2 2 2 2
No. of Actor hidden dims 256 256 256 256 256 256
No. of Critic layers 2 2 2 2 2 2
No. of Critic hidden dims 256 256 256 256 2048 256
Activations in critic network GeLU GeLU Mish ReLU ReLU Mish
Activations in actor network Mish Mish Mish ReLU ReLU Mish
Diffusion steps 5 20 20 20 16 20
Policy delay update 2 2 2 2 2 2
Action gradient steps N/A N/A N/A N/A N/A 20
No. of Gaussian distributions 3 3 N/A N/A N/A N/A
No. of action samples 200 200 N/A N/A N/A N/A
Time-weighted hyperparameter $c$ 0.4 N/A N/A N/A N/A N/A
Time-weighted hyperparameter $d$ -1.8 N/A N/A N/A N/A N/A
Alpha delay update 10,000 10,000 N/A N/A N/A N/A
Noise scale $\lambda$ 0.1 0.1 N/A N/A N/A N/A
Optimizer Adam Adam Adam Adam Adam Adam
Actor learning rate $\cdot 10^{-4} 11$$\cdot 10^{-4}$ $\cdot 10^{-4} 11$$\cdot 10^{-4}$ $3\cdot 10^{-4}$ $1\cdot 10^{-4}$
Critic learning rate $\cdot 10^{-4} 11$$\cdot 10^{-4}$ $\cdot 10^{-4} 11$$\cdot 10^{-4}$ $3\cdot 10^{-4}$ $1\cdot 10^{-4}$
Alpha learning rate $3\cdot 10^{-2}$ $3\cdot 10^{-2}$ N/A N/A $1\cdot 10^{-3}$ N/A
Target entropy ${\rm{dim}}(\mathcal{A})$ ${\rm{dim}}(\mathcal{A})$ N/A N/A $4{\rm{dim}}(\mathcal{A})$ N/A
@@ Appendix E Limitation and Future Work
In this study, we propose the Q-gradient field objective as an auxiliary training loss to provide more informative gradient signals for guiding the diffusion policy. However, algorithms such as PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) do not explicitly learn a Q-function, making it challenging to directly integrate the diffusion policy of DACERv2 and its associated loss function with these methods. This indicates that the generality of our method is currently affected by the presence of value functions. Future work could explore reformulating the auxiliary objective into a purely trajectory-based form, thereby enabling integration with methods that rely solely on policy gradients.
Appendix F Extra ablation study
We conducted an ablation study on the Humanoid-v3 task to examine the effect of normalizing the Q-gradient. The results presented in Fig. 13 demonstrate that normalization method consistently enhance performance returns.
Figure 13: Ablation on the normalization of Q-function.
Appendix G LLM statement
Large Language Models (LLMs) were employed solely for language refinement in this paper. Specifically, we used them to polish grammar, improve clarity, and enhance the academic style of our writing. The role of LLMs was limited to editing and improving the presentation of the text, without contributing to the technical content.